OpenStudy (cutiecomittee123):

How in the world do I solve this? sec^2 (x)+sec (x)-2=0 I know you have to simlify it so I added two on both sides so far and I am stuck after that

OpenStudy (amistre64):

its a disguised quadratic equation

OpenStudy (amistre64):

do you know your quadratic formula?

OpenStudy (cutiecomittee123):

yes

OpenStudy (cutiecomittee123):

but right now I am learning the basics of solving this and I know that you must first simplify

OpenStudy (amistre64):

if you are learning the basics ,,, then complete the square, which is just the proofing of the quad formula

OpenStudy (amistre64):

in this case, id use the quad formula if you are expected to already know it.

OpenStudy (cutiecomittee123):

How would this plug into the quadratic formula?

OpenStudy (amistre64):

let u=sec(x) u^2 +u - 2 = 0 what is your formula?

OpenStudy (cutiecomittee123):

a=1 b=1 c=-2

OpenStudy (amistre64):

good

OpenStudy (cutiecomittee123):

-1+-sqrt 1^2-4(1)(-2) all of that over 2

OpenStudy (cutiecomittee123):

but then you get a negative in the sqrt

OpenStudy (amistre64):

therefore: \[u=sec(x)=\frac{-1\pm\sqrt{1-4(1)(-2)}}{2(1)}\] recall that sec = 1/cos soo cos(x)=2/(1 +- sqrt(9)) x = cos^(-1) (....)

OpenStudy (amistre64):

no, no you dont :)

OpenStudy (amistre64):

1 +- 3 = -2,4 soo cos(x) = -1, 1/2 is our solutions

OpenStudy (cutiecomittee123):

oh yeah i forgot... so it would be the sqrt of 9 which is 3 so -1+3/2 and -1-3/2

OpenStudy (amistre64):

if you go with sec yes,

OpenStudy (cutiecomittee123):

so 1 and -2

OpenStudy (amistre64):

sounds good to me

OpenStudy (amistre64):

arc trig them to get x

OpenStudy (cutiecomittee123):

but the thing is there are three solutions that need to be found and that is only two

OpenStudy (amistre64):

thats not the end of it ... when you inverse the trig you will sovle for x

OpenStudy (amistre64):

\[sec(x)=\frac{1\pm3}{2}\]

OpenStudy (amistre64):

opps, forgot the 1 is negtaive

OpenStudy (cutiecomittee123):

you still get 1, and -2

OpenStudy (amistre64):

\[sec(x)=\frac{-1\pm3}{2}\] \[sec(x)=-2,1\] \[cos(x)=1,-\frac{1}{2}\] cos(x)=-1/2 has 2 values between 0 and 2pi cos(x)=1 only has 1 value

OpenStudy (amistre64):